OpenStudy (phagey):

A painter charges a flat fee of $125 for supplies, $37 per hour of painting and $32 per gallon of paint used. Which of the following is the closet approximation to the total cost of a job that requires 14 hours of work and 12 gallons of paint A.$230 B.$800 C.$930 D.$1,530

OpenStudy (studio-games):

Ok let me see

OpenStudy (studio-games):

1027

OpenStudy (phagey):

so D?

OpenStudy (studio-games):

bloody hell

OpenStudy (phagey):

but how do i do the problem?

OpenStudy (studio-games):

1 second

OpenStudy (studio-games):

oh You do it as it say. Do 37 times 14 and so and so

OpenStudy (phagey):

and 32 times 12?

OpenStudy (nfcfox):

@phagey Yes, add that to what studio said and the flat fee. You will have your answer.

OpenStudy (studio-games):

Yes 384

OpenStudy (studio-games):

then 37 * 14 is 518

OpenStudy (studio-games):

and add 127

OpenStudy (studio-games):

I dont understand though. It isnt showing up as any of the answers???

OpenStudy (studio-games):

oh CLOSEST aproximation...

OpenStudy (anonymous):

well 32 x 12 = 384 + 37 x 10 = 370 + 4 x 37 = 148 so the anwser is ...

OpenStudy (phagey):

so d or c?

OpenStudy (studio-games):

I would say C because it is closer to 930 than 1530.

OpenStudy (studio-games):

Good luck

OpenStudy (phagey):

ok,im confused because everyone is giving me different answers

OpenStudy (studio-games):

C

OpenStudy (studio-games):

There

OpenStudy (studio-games):

Hope that helps you and good luck on your test.

OpenStudy (anonymous):

c

OpenStudy (studio-games):

It does not cost 125 per gallon...

OpenStudy (studio-games):

Read it again.

OpenStudy (studio-games):

It is 32 $ per gallon

OpenStudy (studio-games):

@XxStudyingStuffxX is WAY off

OpenStudy (phagey):

ok so its c

OpenStudy (studio-games):

Do you understand now phagey?

OpenStudy (studio-games):

Yes and good luck!

OpenStudy (phagey):

yes

OpenStudy (studio-games):

Thanks!

OpenStudy (anonymous):

yes phagey

OpenStudy (xxstudyingstuffxx):

Sorry. I didnt see the price for paint. But he wasn't painting with his hands. He probably had a brush and tray or something.

OpenStudy (xxstudyingstuffxx):

@studio-games is right. 32 * 12 = 384. The paint costed $384. ||| 37 * 14 = 518. The work costed $518. ||| C is the answer. Hope it helps!